OpenStudy (anonymous):

PB has endpoints at -2 and 12. What is the coordinate of its midpoint?

OpenStudy (anonymous):

wouldn't it be -2+12=10/2=5 right

OpenStudy (tkhunny):

Please never do that. Notice how you have written -2 + 12 = 5. This is utterly incorrect. \(\dfrac{-2+12}{2} = \dfrac{10}{2} = 5\) "=" means things. Don't make it a liar,